in a reaction 35.0 mL of 0.718 M copper II chloride reacts with 45.0 mL of 0.429 M potassium phosphate. Determine the molarity of each of the spectator ions.


1
Expert's answer
2021-02-22T05:54:11-0500

Copper II chloride (CaCl2) solution (35.0 mL) contains 35.0 mL × 0.718 M = 0.035 L × 0.718 mol/L = 0.025 mol.

From here, the amounts of ions in the initial solution are:

n(Ca2+) = 0.025 mol

n(Cl-) = 0.025 × 2 = 0.05 mol

Potassium phosphate (K3PO4) solution (45.0 mL) contains 45.0 mL × 0.429 M = 0.045 L × 0.429 mol/L = 0.019 mol.

From here, the amounts of ions in the initial solution are:

n(K+) = 0.019 mol × 3 = 0.057 mol

n(PO43-) = 0.019 mol

The final volume of the solution is:

V = 35.0 mL + 45.0 mL = 80.0 mL = 0.080 L

From here, the molarity of each ion in the final solution is:

c(Ca2+) = n(Ca2+) / V = 0.025 mol / 0.080 L = 0.31 M

c(Cl-) = n(Cl-) / V = 0.05 mol / 0.080 L = 0.625 M

c(K+) = n(K+) / V = 0.057 mol / 0.080 L = 0.071 M

c(PO43-) = n(PO43-) / V = 0.019 mol / 0.080 L = 0.24 M